Find the sum of 7th term of the G.P 12, 6, 3
Answer by htmentor(1343)

The n-th term of a geometric progression is written
a_n = a*r^(n-1), where a = the 1st term and r is the common ratio
In this case, a = 12 and r = 6/12 = 1/2
So the n-th term is a_n = 12(1/2)^(n-1)
Thus the 7th term is a_7 = 12(1/2)^6 = 12/64 = 3/16